# Have Faith in Faithfulness: Going Beyond Circuit Overlap When Finding Model Mechanisms

### v1

@@ Abstract
Many recent language model (LM) interpretability studies have adopted the circuits framework, which aims to find the minimal computational subgraph, or circuit, that explains LM behavior on a given task. Most studies determine which edges belong in a LM’s circuit by performing causal interventions on each edge independently, but this scales poorly with model size. Edge attribution patching (EAP), gradient-based approximation to interventions, has emerged as a scalable but imperfect solution to this problem. In this paper, we introduce a new method—EAP with integrated gradients (EAP-IG)—that aims to better maintain a core property of circuits: faithfulness. A circuit is faithful if all model edges outside the circuit can be ablated without changing the model’s performance on the task; faithfulness is what justifies studying circuits, rather than the full model. Our experiments demonstrate that circuits found using EAP are less faithful than those found using EAP-IG, even though both have high node overlap with circuits found previously using causal interventions. We conclude more generally that when using circuits to compare the mechanisms models use to solve tasks, faithfulness, not overlap, is what should be measured.
@@ 1 Introduction
In this paper, we answer these questions, making three main contributions. First, we draw on earlier attribution literature (Sundararajan et al., 2017) to propose a new circuit-finding method—edge attribution patching with integrated gradients (EAP-IG)—that finds more faithful circuits (Figure 1). Second, we test the faithfulness of not only EAP and EAP-IG circuits, but also circuits found via activation patching scores, the ground truth that EAP and EAP-IG scores aim to approximate. We find that EAP-IG circuits are more faithful than EAP circuits on many tasks, though activation patching circuits still sometimes outperform both. Third, we empirically investigate the relationship between overlap and faithfulness, and show the former does not always imply the latter. We will release our code in this repository.
@@ 2 The Circuits Framework and Circuit Finding
Computational graph: A transformer LM’s computational graph is a digraph describing the computations it performs. It flows from the LM’s inputs to the unembedding that projects its activations into vocabulary space. We define this digraph’s nodes to be the LM’s attention heads and MLPs, though other levels of granularity, e.g., neurons, are possible. Edges specify where a node’s output goes; a node $v$’s input is the sum of the outputs of all nodes $u$ with an edge to $v$. A circuit is a subgraph of this that connects the inputs to the logits.
@@
Typically, circuits are localized to a small portion of the LM’s edges; the vast majority of the LM’s activations are hence corrupted. However, if the circuit is good, the model’s behavior on the clean inputs should be roughly the same with or without these interventions. We verify this by computing our metric $M$ on the model with interventions. This should be close to $M$’s value when running the model without interventions on clean inputs.
@@
where $\nabla_{v}L(s)$ is the gradient of $L$ with respect to the input of $v$. Note that EAP estimates the change in loss $L$ caused by intervening, rather than the change in metric $M$, but we can turn a metric into a loss by taking $L(x)=-M(x)$. The change in loss (or “score”) for each edge can be computed with one forward pass on corrupted inputs, and one forward and backward pass on clean inputs. Having scored each edge, one then selects which edges fall in the circuit, e.g. by taking the top-$n$ edges by absolute score. In terms of node overlap, this method recovers manually found circuits with high precision and recall (Syed et al., 2023). [^2] Contemporaneous work (Ferrando & Voita, 2024; Kramár et al., 2024) has also introduced new methods for efficient circuit finding, though not using the lens of faithfulness. Due to their contemporaneity (and lack of publicly available implementation) we are unable to compare with them.
@@ 3 EAP with Integrated Gradients
where $m$ is the number of steps used to approximate the integral, $z$ is a sequence of token embeddings for one input, and $z^{\prime}$ is the token embeddings of the distinct, baseline input. [^3] Previously, we ran $M$ on string inputs ($s$ or $s^{\prime}$); here, we specify the inputs as token embeddings, as IG runs our model on a blend of two inputs, which is possible only in embedding (not string) space. We can then combine EAP and IG into a new method, edge attribution patching with integrated gradients (EAP-IG). The EAP-IG score of an edge $(u,v)$ is
@@ 4 Evaluating Edge Attribution Faithfulness
How can we tell which of our two circuit finding techniques is better? Unlike earlier work, we forgo component overlap measures [^4] EAP-IG does still successfully capture component overlap, though; see Appendix B in favor of comparing the faithfulness of the circuits found by EAP and EAP-IG, across six different tasks. We also compare these to circuits based on scores from activation patching; recall that it is precisely these scores that EAP approximates. Computing these values is slow, but the faithfulness of these circuits represents the performance of EAP and EAP-IG with zero approximation error. Note, however, that activation patching circuits are imperfect; scoring each edge independently and then finding a circuit may be less accurate than e.g. ACDC.
@@ 4.1 Tasks
Country–Capital: In the Country-Capital task, models receive input like “Tirana, the capital of” and must output the corresponding country (Albania). Corrupted instances contain another capital (e.g. Brasilia) instead. We measure performance using logit diff: the logit of the correct country (Albania) minus the logit of the corrupted country (Brazil). We create a Country–Capital dataset from the world’s countries and capitals, with 190 examples.
@@ 4.2 Methods
For each task, we score each model edge using EAP, EAP-IG, and activation patching. We convert each task metric $M$ (higher is better) to a loss $L$ (lower is better) by multiplying $M$ by -1; we also run EAP-IG and activation patching with KL divergence as the loss. Then, for $n=30,40,\ldots,100,200,\ldots,1000$, we select a circuit of $n$ edges using a greedy search procedure; if no circuit is faithful at $n=1000$, we consider up to $n=2000$. We use greedy search as it produces more faithful circuits than sorting the edges by absolute score [^5] In both the greedy and top-$n$ cases, we select edges based on absolute scores. Selecting edges with the highest (non-absolute) scores would yield a better-performing circuit, but it would also discard components that are significant due to their negative effect on model performance. and taking the top $n$ edges, as in Syed et al. (2023). For the full greedy algorithm, and details on finding circuits using scored edges, see Appendix D. In general, larger $n$ should yield circuits that are more faithful but less localized and interpretable. We aim to find circuits that are small (containing 1-2% of edges) and faithful (recovering $\geq$85of model performance).
Having found a circuit, we (recursively) prune all nodes and edges that have no children, or no parents. A node with no parent edges in the circuit would have all inputs corrupted, equivalent to not being in the circuit at all; childless nodes can likewise be discarded. Finally, we compute the circuit’s performance, intervening on the model as described in Section 2, and applying the appropriate task metric. Because each task’s metric has a different range, we create a normalized faithfulness score to enable cross-task comparison. Let $b$ and $b^{\prime}$ be the whole model’s performance on the clean and corrupted inputs of the given task respectively; we normalize the circuit’s performance $m$ as $(m-b^{\prime})/(b-b^{\prime})$.
@@ 4.3 Results
In Greater-Than and Country-Capital, there are clearer gaps between EAP-IG and EAP performance, and activation patching does not clearly dominate. In Greater-Than, the gap is small but significant: around 0.1 at the small $n$ that are most interesting for circuit finding. In Country-Capital, this difference is larger ($\geq 0.2$) at all values of $n$.
@@ 4.4 Why Does EAP-IG Work?
Using the activation patching edge scores we obtained in our previous experiment, we investigate why EAP-IG often outperforms EAP. One might hypothesize that, in cases where EAP-IG outperforms EAP, it estimates activation patching scores better than EAP does. However, EAP-IG does not consistently have lower estimation error than EAP, and the difference between EAP and EAP-IG’s total estimation errors is dwarfed by the scale of the errors themselves; see Appendix E for details EAP-IG scores are, however, better correlated with activation patching scores than EAP scores are. Figure 4 (left) shows that the Pearson correlation between activation patching scores and EAP-IG scores is consistently high (above 0.8), while the correlation of activation patching and EAP is always lower, though all correlations are significant ($p<0.01$).
@@
Despite this, EAP-IG does not better approximate the edge ranking induced by activation patching scores. Consider the ranking induced on edges by absolute activation patching, EAP, and EAP-IG scores. Activation patching’s ranking has a significantly higher Kendall correlation (Appendix E) with EAP’s ranking than EAP-IG’s ranking. However, this measure is flawed: it considers ranking similarity across all edges, but most edges are near irrelevant for circuits. What matters is how EAP and EAP-IG rank the most important edges.
@@ 5 Overlap and Faithfulness
Our previous experiments show that circuits found with EAP are not always faithful, despite their high node overlap with manually found circuits. This evidence (see Appendix F for more) pertains to within-task relationship of overlap and faithfulness, circuit but overlap can also be studied in cross-task scenarios (Merullo et al., 2024). So, we ask: to what extent does circuit overlap capture the similarity of mechanisms that models use across tasks?
To answer this, we analyze cross-task circuit overlap and faithfulness, i.e. the overlap and faithfulness obtained when comparing and running circuits across tasks. For each task, we take the smallest circuit we have found that achieves at least 85% faithfulness on the task; this entails using a circuit based on scores from real activation-patching or EAP-IG-KL. We then compute two overlap metrics: the intersection over union (IoU, also called Jaccard similarity) between the circuits’ nodes, and the IoU between their edges. We also compute the statistical significance of each overlap; see Appendix G for details.
We then compute cross-task faithfulness: for every pair of circuits $C_{1},C_{2}$, we run $C_{1}$ on $C_{2}$’s task (and vice-versa), recording the faithfulness achieved. We do under the assumption that if the model uses the same mechanisms to solve both tasks, the tasks’ circuits should be faithful to one another. However, there is no guarantee that any pair of circuits will be faithful across our current set of tasks. This poses a problem: we cannot assess whether overlap predicts faithfulness, without any task pairs whose circuits exhibit cross-task faithfulness. We resolve this issue by studying two additional tasks: Greater-Than (Price) and Greater-Than (Sequence). These are simple variants on the Greater-Than task; they consist of inputs like “The ring’s price ranges from \$ 1741 to \$17” and “1629, 1671, 1692, 1741, 17”, respectively. Like Greater-Than, their corrupted inputs include a $01$, and they use prob diff; for more details, see Appendix H. We include these tasks because Hanna et al. (2023) report that the Greater-Than circuit is both similar (though not identical) to these tasks’ circuits, and generally faithful when run on them.
@@
The cross-tasks relationships suggested by each of our experiments are broadly similar, with small but key differences. Figure 5 (left) indicates that while all task pairs have at least a small level of node overlap (0.25 IoU), certain pairs have notably higher overlap. As expected, the Greater-Than task variants all have high overlap ($\geq$0.6) with one other. However, the Hypernymy and Country-Capital tasks also have high node IoU. This likely stems from the similar nature of the two tasks; in both, the LM must retrieve a property of a given word. In the former, the LM retrieves the word’s hypernym; in the latter, the LM retrieves the word’s (city’s) location. In contrast, the baseline edge overlap (Figure 5, center) between tasks is very small, often between 0.05 and 0.15 IoU. As in the node overlap scenario, the Greater-Than task variants have high edge overlap compared to other task pairs (0.4 IoU). However, this is not true of the Hypernymy and Country-Capital tasks.
@@
The pairwise faithfulness of tasks (Figure 5, right) is more complicated. As in the prior two scenarios, the Greater-Than task variants have high cross-task faithfulness. However, while the Hypernymy circuit ($y$-axis) is highly faithful on the Country-Capital task ($x$-axis), the reverse is not true. This is possible because faithfulness, unlike the other measures, which are based on IoU, is asymmetrical. As the hypernymy circuit is much larger than other circuits, it is likely better able to be faithful to a variety of tasks: it has moderate faithfulness on SVA and Gender-Bias as well. Most other task pairs have near-zero faithfulness.
@@
This raises the question: to what extent does overlap inform faithfulness? The two are indeed strongly related: Pearson correlation of faithfulness with node overlap 0.87, and its correlation with edge overlap is 0.86; both are highly significant ($p<0.001$). This is expected—component overlap must have some positive impact on faithfulness. A linear regression from node overlap to faithfulness achieves an $R^{2}$ of $0.75$, while using edge overlap achieves $0.734$, and combining the two yields $R^{2}=0.754$; these are all strong fits.
@@
Despite this, we caution against concluding that overlap is a good way to predict faithfulness. The most interesting points are those where task overlap is moderate; it is unsurprising that circuits with near zero overlap are unfaithful across tasks, and that circuits with perfect overlap are. But when overlap is moderate, it does not predict faithfulness. The node overlap between Country-Capital and Hypernymy is 0.58, but the Hypernymy circuit achieves a faithfulness of 0.61 on Country-Capital; in the reverse case, faithfulness is 0.2. For one value of overlap, we have two very different values of faithfulness; no linear regression based on overlap can capture this. Even across tasks, there are challenges: the node overlap between Greater-Than (Price) and Greater-Than (Sequence) is 0.64, much like that of Hypernymy and Country-Capital, but cross-task faithfulness is high either way. We thus conclude that overlap is not a good predictor of cross-task faithfulness when overlaps are moderate.

### v2

@@ Abstract
Many recent language model (LM) interpretability studies have adopted the circuits framework, which aims to find the minimal computational subgraph, or circuit, that explains LM behavior on a given task. Most studies determine which edges belong in a LM’s circuit for a task by performing causal interventions on each edge independently, but this scales poorly with model size. As a solution, recent work has proposed edge attribution patching (EAP), a scalable gradient-based approximation to interventions. In this paper, we introduce a new method—EAP with integrated gradients (EAP-IG)—that aims to efficiently find circuits while better maintaining one of their core properties: faithfulness. A circuit is faithful if all model edges outside the circuit can be ablated without changing the model’s behavior on the task; faithfulness is what justifies studying circuits, rather than the full model. Our experiments demonstrate that circuits found using EAP-IG are more faithful than those found using EAP, even though both have high node overlap with reference circuits found using causal interventions. We conclude more generally that when comparing circuits, measuring overlap is no substitute for measuring faithfulness.
@@ 1 Introduction
In this paper, we answer these questions, making three main contributions. First, we draw on earlier attribution literature (Sundararajan et al., 2017) to propose a new circuit-finding method—edge attribution patching with integrated gradients (EAP-IG)—that finds more faithful circuits (Figure 1). Second, we test the faithfulness of not only EAP and EAP-IG circuits, but also circuits found via activation patching scores, the ground truth that EAP and EAP-IG scores aim to approximate. We find that EAP-IG circuits are more faithful than EAP circuits on many tasks, though activation patching circuits still sometimes outperform both. Third, we empirically investigate the relationship between overlap and faithfulness, and show the former does not always imply the latter. We make the code for this paper’s experiments available at https://github.com/hannamw/eap-ig-faithfulness; we also release an implementation of EAP-IG and other circuit finding methods at https://github.com/hannamw/eap-ig.
@@ 2 The Circuits Framework and Circuit Finding
Computational graph: A transformer LM’s computational graph is a digraph describing the computations it performs. It flows from the LM’s inputs to the unembedding that projects its activations into vocabulary space. We define this digraph’s nodes to be the LM’s attention heads and MLPs, though other levels of granularity, e.g., neurons, are possible. Edges specify where a node’s output goes; a node $v$’s input is the sum of the outputs of all nodes $u$ with an edge to $v$ [^1] In modern transformer LM architectures, a node’s input is the sum of the outputs of all previous nodes; for details, see Elhage et al. (2021). Thus, any given node has edges to all downstream nodes. A circuit is a subgraph of this that connects the inputs to the logits.
@@
Typically, circuits are localized to a small portion of the LM’s edges; the vast majority of the LM’s activations are hence corrupted. However, if the circuit is correct, the model’s behavior on the clean inputs should be roughly the same with or without these interventions. We verify this by computing our metric $M$ on the model with interventions. This should be close to $M$’s value when running the model without interventions on clean inputs.
We note that our particular test of faithfulness, which employs activation patching using corrupted examples, is not universal; others have used mean ablations over a corrupted dataset (Prakash et al., 2024), or even zero ablations (Olsson et al., 2022). Chan et al. (2022) advise against zero ablations, which may fall outside models’ normal activation distribution, causing models to break even when unimportant nodes are ablated. Though they also advise mean ablations for the same reason, these seem in practice to be a reasonable alternative to patching, which we believe would yield similar results.
@@
where $\nabla_{v}L(s)$ is the gradient of $L$ with respect to the input of $v$. Note that EAP estimates the change in loss $L$ caused by intervening, rather than the change in metric $M$, but we can turn a metric into a loss by taking $L(x)=-M(x)$. The change in loss (or “score”) for each edge can be computed with one forward pass on corrupted inputs, and one forward and backward pass on clean inputs. Having scored each edge, one then selects which edges fall in the circuit, e.g. by taking the top-$n$ edges by absolute score. In terms of node overlap, this method recovers manually found circuits with high precision and recall (Syed et al., 2023). [^3] While we focus mainly on EAP, other recent work (Ferrando & Voita, 2024; Kramár et al., 2024) has introduced new methods for efficient circuit finding, though not using the lens of faithfulness. Due to their contemporaneity (and lack of publicly available implementation) we are unable to compare with them, but such comparisons would be valuable.
@@ 3 EAP with Integrated Gradients
where $m$ is the number of steps used to approximate the integral, $z$ is a sequence of token embeddings for one input, and $z^{\prime}$ is the token embeddings of the distinct, baseline input. [^4] Previously, we ran $M$ on string inputs ($s$ or $s^{\prime}$); here, we specify the inputs as token embeddings, as IG runs our model on a blend of two inputs, which is possible only in embedding (not string) space. We can then combine EAP and IG into a new method, edge attribution patching with integrated gradients (EAP-IG). [^5] Contemporaneous work (Marks et al., 2024) has proposed a method similar to EAP-IG, which interpolates between node outputs, rather than input embeddings. This method is slower, as its cost scales with model depth, and achieves performance similar to that of our EAP-IG. See Appendix D for an in-depth comparison of the two methods and additional baselines. The EAP-IG score of an edge $(u,v)$ is
@@ 4 Evaluating Edge Attribution Faithfulness
How can we tell which of our two circuit finding techniques is better? Unlike earlier work, we forgo component overlap measures [^6] EAP-IG does still successfully capture component overlap, though; see Appendix B in favor of comparing the faithfulness of the circuits found by EAP and EAP-IG, across six different tasks, on GPT-2 small (Radford et al., 2019). We also compare these to circuits based on scores from activation patching; recall that it is precisely these scores that EAP approximates. Computing these values is slow, but the faithfulness of these circuits represents the performance of EAP and EAP-IG with zero approximation error. Note, however, that activation patching circuits are imperfect; scoring each edge independently and then finding a circuit may be less accurate than e.g. ACDC.
@@ 4.1 Tasks
Capital–Country: In the Capital-Country task, models receive input like “Tirana, the capital of” and must output the corresponding country (Albania). Corrupted instances contain another capital (e.g. Brasilia) instead. We measure performance using logit diff: the logit of the correct country (Albania) minus the logit of the corrupted country (Brazil). We create a Capital-Country dataset from the world’s countries and capitals, with 190 examples.
@@ 4.2 Methods
For each task, we score each model edge using EAP, EAP-IG, and activation patching. We convert each task metric $M$ (higher is better) to a loss $L$ (lower is better) by multiplying $M$ by -1; we also run EAP-IG and activation patching with KL divergence as the loss. Then, for $n=30,40,\ldots,100,200,\ldots,1000$, we select a circuit of $n$ edges using a greedy search procedure; if no circuit is faithful at $n=1000$, we consider up to $n=2000$. We use greedy search as it produces more faithful circuits than sorting the edges by absolute score [^7] In both the greedy and top-$n$ cases, we select edges based on absolute scores. Selecting edges with the highest (non-absolute) scores would yield a better-performing circuit, but it would also discard components that are significant due to their negative effect on model performance. and taking the top $n$ edges, as in Syed et al. (2023). For the full greedy algorithm, and details on finding circuits using scored edges, see Appendix E. In general, larger $n$ should yield circuits that are more faithful but less localized and interpretable. We aim to find circuits that are small (containing 1-2% of edges) and faithful (recovering $\geq$85of model performance).
Having found a circuit, we (recursively) prune all nodes and edges that have no children, or no parents. A node with no parent edges in the circuit would have all inputs corrupted, equivalent to not being in the circuit at all; childless nodes can likewise be discarded. Finally, we compute the circuit’s performance, intervening on the model as described in Section 2, and applying the appropriate task metric. Because each task’s metric has a different range, we create a normalized faithfulness score to enable cross-task comparison. Let $b$ and $b^{\prime}$ be the whole model’s performance on the clean and corrupted inputs of the given task respectively; we normalize the circuit’s performance $m$ as $(m-b^{\prime})/(b-b^{\prime})$. We implement these and all following experiments in the TransformerLens library (Nanda & Bloom, 2022).
@@ 4.3 Results
In Greater-Than and Capital-Country, there are clearer gaps between EAP-IG and EAP performance, and activation patching does not clearly dominate. In Greater-Than, the gap is small but significant: around 0.1 at the small $n$ that are most interesting for circuit finding. In Capital-Country, this difference is larger ($\geq 0.2$) at all values of $n$.
@@ 4.4 Why Does EAP-IG Work?
Using the activation patching edge scores we obtained in our previous experiment, we investigate why EAP-IG often outperforms EAP. One might hypothesize that, in cases where EAP-IG outperforms EAP, it estimates activation patching scores better than EAP does. However, EAP-IG does not consistently have lower estimation error than EAP, and the difference between EAP and EAP-IG’s total estimation errors is dwarfed by the scale of the errors themselves; see Appendix F for details EAP-IG scores are, however, better correlated with activation patching scores than EAP scores are. Figure 4 (left) shows that the Pearson correlation between activation patching scores and EAP-IG scores is consistently high (above 0.8), while the correlation of activation patching and EAP is always lower, though all correlations are significant ($p<0.01$).
@@
Despite this, EAP-IG does not better approximate the edge ranking induced by activation patching scores. Consider the ranking induced on edges by absolute activation patching, EAP, and EAP-IG scores. Activation patching’s ranking has a significantly higher Kendall correlation (Appendix F) with EAP’s ranking than EAP-IG’s ranking. However, this measure is flawed: it considers ranking similarity across all edges, but most edges are near irrelevant for circuits. What matters is how EAP and EAP-IG rank the most important edges.
@@ 5 Overlap and Faithfulness
Our previous experiments show that circuits found with EAP are not always faithful, despite their high node overlap with manually found circuits. This evidence (see Appendix G for more) pertains to within-task relationship of overlap and faithfulness, circuit but overlap can also be studied in cross-task scenarios (Merullo et al., 2024). So, we ask: to what extent does circuit overlap capture the similarity of mechanisms that models use across tasks?
To answer this, we analyze cross-task circuit overlap and faithfulness, i.e. the overlap and faithfulness obtained when comparing and running circuits across tasks. For each task, we take the smallest circuit we have found that achieves at least 85% faithfulness on the task; this entails using a circuit based on scores from real activation-patching or EAP-IG-KL. We then compute two overlap metrics: the intersection over union (IoU, also called Jaccard similarity) between the circuits’ nodes, and the IoU between their edges. We also compute the statistical significance of each overlap; see Appendix H for details.
We then compute cross-task faithfulness: for every pair of circuits $C_{1},C_{2}$, we run $C_{1}$ on $C_{2}$’s task (and vice-versa), recording the faithfulness achieved. We do under the assumption that if the model uses the same mechanisms to solve both tasks, the tasks’ circuits should be faithful to one another. However, there is no guarantee that any pair of circuits will be faithful across our current set of tasks. This poses a problem: we cannot assess whether overlap predicts faithfulness, without any task pairs whose circuits exhibit cross-task faithfulness.
We resolve this issue by studying three additional tasks: Greater-Than (Price), Greater-Than (Sequence), and Country-Capital. The first two are simple variants on the Greater-Than task; they consist of inputs like “The ring’s price ranges from \$ 1741 to \$17” and “1629, 1671, 1692, 1741, 17”, respectively. Like Greater-Than, their corrupted inputs include a $01$, and they use prob diff; for more details, see Appendix I. We include these tasks because Hanna et al. (2023) report that the Greater-Than circuit is both similar (though not identical) to these tasks’ circuits, and generally faithful when run on them. The last task, Country-Capital, is the reverse of the Capital-Country task, consisting of examples like “Albania, whose capital”, where the desired answer is “Tirana”. We include this task because we hypothesize its circuit may overlap with that of Capital-Country.
@@
The cross-tasks relationships suggested by each of our experiments are broadly similar, with small but key differences. Figure 5 (left) indicates that while all task pairs have at least a small level of node overlap (0.25 IoU), certain pairs have notably higher overlap. As expected, the Greater-Than task variants all have high overlap ($\geq$0.6) with one other, as do the Capital-Country and Country-Capital tasks. However, the Hypernymy task also has high node IoU with the latter two tasks. This likely stems from the similar nature of the tasks; in both, the LM must retrieve a property of a given word. In Hypernymy, the LM retrieves the word’s hypernym; in the latter pair, the LM retrieves the word’s (city’s) location or the word’s (country’s) capital.
In contrast with node overlap, the baseline edge overlap (Figure 5, center) between tasks is very small, often between 0.05 and 0.15 IoU. As in the node overlap scenario, the Greater-Than task variants, as well as Country-Capital and Capital-Country, have high edge overlap compared to other task pairs (0.4 IoU). However, this is not true of the Hypernymy and Capital-Country or Country-Capital tasks.
@@
The pairwise faithfulness of tasks (Figure 5, right) is more complicated. As in the prior two scenarios, the Greater-Than task variants have high cross-task faithfulness, as do Capital-Country and Country-Capital [^9] This is an interesting result in light of the reversal curse: Berglund et al. (2024) find that LMs trained on “A is B” do not learn “B is A”. Here, where the model has likely been trained on both facts, it uses similar mechanisms to retrieve both.. However, while the Hypernymy circuit ($y$-axis) is highly faithful on the Capital-Country and Country-Capital tasks ($x$-axis), the reverse is not true. This is possible because faithfulness, unlike the other measures, which are based on IoU, is asymmetrical. As the hypernymy circuit is much larger than other circuits, it is likely better able to be faithful to a variety of tasks: it has moderate faithfulness on SVA and Gender-Bias as well. Most other task pairs have near-zero faithfulness.
@@
This raises the question: to what extent does overlap inform faithfulness? The two are indeed strongly related: Pearson correlation of faithfulness with node overlap 0.86, and its correlation with edge overlap is 0.83; both are highly significant ($p<0.001$). This is expected—component overlap must have some positive impact on faithfulness. A linear regression from node overlap to faithfulness achieves an $R^{2}$ of $0.74$, while using edge overlap achieves $0.70$, and combining the two yields $R^{2}=0.74$; these are all strong fits.
@@
Despite this, we caution against concluding that overlap is a good way to predict faithfulness. The most interesting points are those where task overlap is moderate; it is unsurprising that circuits with near zero overlap are unfaithful across tasks, and that circuits with perfect overlap are. But when overlap is moderate, it does not predict faithfulness. The node overlap between Capital-Country and Hypernymy is 0.58, but the Hypernymy circuit achieves a faithfulness of 0.61 on Capital-Country; in the reverse case, faithfulness is 0.2. For one value of overlap, we have two very different values of faithfulness; no linear regression based on overlap can capture this. Even across tasks, there are challenges: the node overlap between Greater-Than (Price) and Greater-Than (Sequence) is 0.64, much like that of Hypernymy and Capital-Country, but cross-task faithfulness is high either way. We thus conclude that overlap is not a good predictor of cross-task faithfulness when overlaps are moderate.